# No Language is an Island: Unifying Chinese and English in Financial Large Language Models, Instruction Data, and Benchmarks

### v1

@@ Abstract
While the progression of Large Language Models (LLMs) has notably propelled financial analysis, their application has largely been confined to singular language realms, leaving untapped the potential of bilingual Chinese-English capacity. To bridge this chasm, we introduce ICE-PIXIU [^1] https://github.com/The-FinAI/PIXIU, seamlessly amalgamating the ICE-INTENT model and ICE-FLARE benchmark for bilingual financial analysis. ICE-PIXIU uniquely integrates a spectrum of Chinese tasks, alongside translated and original English datasets, enriching the breadth and depth of bilingual financial modeling. It provides unrestricted access to diverse model variants, a substantial compilation of diverse cross-lingual and multi-modal instruction data, and an evaluation benchmark with expert annotations, comprising 10 NLP tasks, 20 bilingual specific tasks, totaling 1,185k datasets. Our thorough evaluation emphasizes the advantages of incorporating these bilingual datasets, especially in translation tasks and utilizing original English data, enhancing both linguistic flexibility and analytical acuity in financial contexts. Notably, ICE-INTENT distinguishes itself by showcasing significant enhancements over conventional LLMs and existing financial LLMs in bilingual milieus, underscoring the profound impact of robust bilingual data on the accuracy and efficacy of financial NLP.
@@ 1 Introduction
Beyond the efforts in English-centric models, Chinese financial LLMs (FinLLMs) have also made notable advances. XuanYuan2.0 (Zhang, 2023) exemplifies this trend, focusing on the intricacies of Chinese financial language. This has been followed by models such as DISC-FinLLM (Chen et al., 2023) and CFGPT (Li et al., 2023a), which further the capabilities in Chinese financial analysis and reasoning. Recently, PanGu-π (Wang et al., 2023) enhances a pretrained model for better performance in Chinese financial examinations. Moreover, certain business-oriented LLMs like TigerBot and LightGPT remain closed-source (Wang, 2023).
@@
BloombegGPT (Wu et al., 2023) 50B en 0 5 0 5 0 $usym2717 \usym$2714 03/30/23
InvestLM (Yang et al., 2023c) 65B en 0 9 0 9 0 $usym2717 \usym$2714 09/14/23
FinGPT (Wu et al., 2023) 7/13B en 0 6 0 6 0 $usym2717 \usym$2714 11/10/23
XuanYuan2.0 (Zhang, 2023) 176B zh - 0 0 1 0 $usym2717 \usym$2714 05/19/23
FinMA (Xie et al., 2023b) 7/13B en 0 8 0 8 0 $usym2714 \usym$2717 06/01/23
[…]
ICE-INTENT 7B zh, en 12 6 12 6 3 $usym2714 \usym$2717 12/10/23
[…]
@@
In the creation of ICE-INTENT, we meticulously gathered 40 datasets, which consisted of 1,185,076 raw data, 603,940 instruction fine-tuning data, and 95,091 evaluation data, strategically covering a wide range of financial tasks. This ensemble, as detailed in Table 1, comprises 18 specific task on Chinese-English bilingual instruction datasets. This diverse array not only underscores our model’s commitment to linguistic breadth but also to task-specific depth, ensuring ICE-INTENT’s proficiency across various financial scenarios  [^2] Detailed examples are shown in Appendix Figure 7-8. Complementing this, ICE-FLARE, our cross-lingual evaluation benchmark, incorporates 21 evaluation specific tasks—12 in Chinese and 9 in English. Particularly, we innovatively introduced data-out-of-field specific task into ICE-FLARE. This robust framework highlights the unique strength of our approach in providing comprehensive cross-lingual consistency and generalization assessments, crucial for bilingual FinLLMs.
@@ 2 Related Work
Financial Dataset Resource: Financial pre-trained language models (FinPLMs) previously focused on single-task data but now most FinLLMs prioritize specific-task data collection. For instance, BloombergGPT (Wu et al., 2023) integrates general and financial text for training large 500B LLMs. FinLLMs like FinGPT (Yang et al., 2023b; Zhang et al., 2023a) curate English instruction datasets for classification, extraction, and generation NLP tasks. PIXIU  (Xie et al., 2023b) provide open-source, multi-task and multi-modal data covering diverse financial scenarios. BBT-Fin (Lu et al., 2023) Corpus aggregates Chinese financial data including corporate announcements, research reports, social media, and more, while XuanYuan (轩辕) 2.0 (Zhang, 2023) offers a hybrid-tuning dataset. Cornucopia’s (Yu, 2023) 26M instruction dataset covers application areas like insurance and stocks. Most fine-tuning datasets are tailored for monolingual tasks, limiting coverage and open sourcing.
@@ 3.2.1 Chinese Financial Datasets
For ZH-CLS, our dataset covers essential tasks such as sentiment analysis (FinSA), semantic matching (FinSM), news classification (FinNC), negative judgment (FinNJ), answer selection (FinAS) and relationship extraction (FinRE). 1) Sentiment Analysis. It is the process of analyzing and labeling the sentiment expressed in financial texts as "Negative(消极)","Neutral(中性)" and "Negative(消极)" (Sohangir et al., 2018). We used the FE (Lu et al., 2023) dataset sourced from Stockbar and Xueqiu websites. We also utilized the StockB   [^3] https://huggingface.co/datasets/kuroneko5943/stock11 dataset, which focuses on the finance in among data from ten domains. 2) Semantic Matching. It determines if financial texts or sentences are semantically similar ("是(YES)" or "否(NO)"), using two datasets: the Bank Question Corpus (BQC)  (Chen et al., 2018) from Chinese bank customer logs and the Ant Financial Question Matching Corpus (AFQMC)  (Xu et al., 2020) from the Alipay competition, both utilizing binary classification for match identification. 3) Negative Judgment. It focuses on identifying finance and economics-related negativity, distinct from general financial sentiment analysis. Leveraging the FinNSP dataset (Lu et al., 2023) sourced from social media, it uses "是(YES)" to label data containing negative information and "否(NO)" for those without. 4) News Classification. It is to categorize news into 2 or multiple-class label, such as "股票市场(stock market)", “外汇(foreign exchange)", "宏观经济(macroeconomics), and more. We use the FinNL  (Lu et al., 2023) dataset, sourced from financial news websites like ina Finance, Tencent Finance, Phoenix Finance, 36Kr, and Huxiu. It includes binary (NL) and multi-class (NL2) labeled data. 5) Answer Selection. FinEval  (Zhang et al., 2023b) and and AntGroup evaluates financial reasoning by selecting the correct option ("A", "B", "C", "D) in a financial context. The dataset   [^4] https://github.com/alipay/financial_evaluation_dataset includes 4,661 questions covering finance, economy, accounting, and certifications. We’ve compiled data for ten key finance subjects as FinEvalF. 6) Relationship Extraction. It aims to provide the types of connections, dependencies, and associations between entities. We gathered the FinRE dataset (Lu et al., 2023) and named it RE, which includes financial news articles along with head-tail entity pairs, covering 44 relationship categories (e.g., "合作(cooperation)", "持股(shareholding)", etc.), including the "未知(unknown)" relation.
@@ 3.2.2 English Financial Datasets
For EN-EXT, we include the Named Entity Recognition task. We use the CoNLL-2003 dataset  (Alvarado et al., 2015) named as NER including sentences from public financial agreements via U.S. SEC filings. It includes four annotated entity types: Location ("LOC"), Organisation ("ORG"), Person ("PER"), and Miscellaneous ("MISC").
@@
Additionally, we introduce EN-OFT including NER, text summarization (FinTS) and hawkish-dovish classification (FinDC) task. 1) Named Entity Recognition. In NER task, we utilize the FINER-ORD  (Alvarado et al., 2015) dataset, which provides labels for entities such as "PER", "LOC", and "ORG". 2) Text Summarization. For this task, we use two datasets: ECTSUM (Mukherjee et al., 2022) and EDTSUM (Zhou et al., 2021). ECTSUM involves extracting key sentences from texts, while EDTSUM focuses on generating fitting news headlines through abstractive summarization. 3) Hawkish-dovish Classification. FinDC aims to classify sentences from monetary policy texts into a "hawkish" or "dovish" stance, unlike FinSA, using the FOMC  (Shah et al., 2023) dataset.
@@ 4 ICE-FIND: Bilingual Financial Instruction Data
Base on the different task of raw datasets, we further construct our bIlingual Chinese-English Financial InstructioN Dataset (ICE-FIND), covering 13 tasks and 36 datasets with 604k data samples in total, as detailed in Appendix Table 6. This dataset compilation combined raw data with expertly devised prompts, generating 30-40 unique prompts for each dataset. However, it is important to note that expert-authored prompts may not always accurately reflect the quality of the generated results. We assessed these prompts through human evaluations  [^6] Examples and thier translations are shown in Appendix Figure 7-8 focusing on accuracy, naturalness, and informativeness using a 1-3 scale, to identify the most effective ones. These prompts were tested on various LLMs using the ChatALL   [^7] https://github.com/sunner/ChatALL platform, showing satisfactory performance in most cases.
To handle prompt variations across LLMs, we adopted two strategies. For English tasks, we matched each data with all prompts except for FinQA and ConvFinQA. For Chinese tasks, we randomly selected a prompt for each data due to the larger scale of the raw data. Finally, we converted the raw datasets into instruction-tuning samples by collecting human-designed instructions with prompts   [^8] Examples are presented in Appendix Table 9, input texts, and responses. The instruction data was structured in JSON format (Figure 3). {data_id} represents the ID of each data, {prompt} is the task-specific prompt, {conext} refers to the input data, {Your answer} denotes the expected answer, and {Your label} represents the truly answer.
@@ 5 ICE-INTERN: Bilingual Financial Large Language Model
Upon evaluating performance of various backbone LLMs on ICE-FLAR, we explore InternLM-7B to bulid ICE-INTERN-7B with the BiFinID instruction dataset. To assess the impact of different data types on fine-tuning ICE-INTERN, we gradually incorporate the DLC, DLE, DTT, DTE datasets during the fine-tuning process to create variants of ICE-INTERN named ICE-INTERN-dlc-7B, ICE-INTERN-dle-7B, ICE-INTERN-dtt-7B, and ICE-INTERN-full-7B, respectively. We use the same set of fine-tuning hyper-parameters for all models, which are described below. To improve the model’s capability in handling long text, we employ an advanced fine-tuning technique called QLoRA (Hu et al., 2021) to enhance the training process, leading to optimized memory usage and faster training times. Specifically, the training sequences are uniformly cut to a length of 2048 tokens. Then, we use AdamW optimizer with an initial learning rate 5e-5 on weight decay 1e-5 for all parameters. Additionally, the warmup steps are set to 1% of the total training steps. Finally, we fine-tune model total 1 epoch in batch size of 24 on 8 A100 40GB GPUs. Due to the growing volume of fine-tuning data and constraints on our computational resources, we have opted not to proceed with additional fine-tuning of ICE-INTERN using larger parameters.
6 ICE-FLARE: Bilingual Financial Evaluation Benchmark
Based on ICE-FIND, we design our bilingual multi-task evaluation benchmark, ICE-FLARE. The existing benchmarks ([1] FLARE (Xie et al., 2023b), [2] CFBenchmark (Lei et al., 2023), [3] FinCUGE (Chen et al., 2023), [4] FLUE (Shah et al., 2022), [5] Fineval (Zhang et al., 2023b), [6] CGCE (Zhang et al., 2023c)) compared for ICE-FLARE are listed in Table 2. Unfortunately, they are all designed exclusively for either Chinese or English. ICE-FLARE has highlighted its superiority over FLUE in English through FLARE  (Xie et al., 2023b). From Table 2, it is evident that ICE-FLARE introduce bilingual capabilities, significantly surpasses these benchmarks on task diversity and data scalability. Despite FinCUGE having a higher number of financial evaluation tasks, the difference in dataset size compared to ICE-FLARE is nearly 8 times. These public benchmarks have provided valuable references and inspiration for constructing the test set, making ICE-FLARE more comprehensive in terms of task coverage, metric diversity, and sample quantity. However, they still face challenges in evaluating the bilingual capabilities in Chinese-English and severely lack DOF tasks (such as FinDC) to test the robustness of the models. In conclusion, ICE-FLARE, which offers a bilingual evaluation benchmark and a more comprehensive assessment of diversity tasks, is necessary for improving performance and linguistic capabilities of bilingual FinLLMs.
@@ 6 ICE-FLARE: Bilingual Financial Evaluation Benchmark
[…]
@@
[…]
@@
[…]
@@
F1 Accuracy zh NSP 500 $usym2717 \usym$2714 $usym2717 \usym$2717
@@
[…]
@@
F1 Accuracy zh RE 1,489 $usym2717 \usym$2714 $usym2717 \usym$2717
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Open Source $usym2714 \usym$$\usym2714$ $usym2714 \usym$2717
@@
[…]

### v2

@@ Abstract
While the progression of Large Language Models (LLMs) has notably propelled financial analysis, their application has largely been confined to singular language realms, leaving untapped the potential of bilingual Chinese-English capacity. To bridge this chasm, we introduce ICE-PIXIU [^1] https://github.com/YY0649/ICE-PIXIU, seamlessly amalgamating the ICE-INTENT model and ICE-FLARE benchmark for bilingual financial analysis. ICE-PIXIU uniquely integrates a spectrum of Chinese tasks, alongside translated and original English datasets, enriching the breadth and depth of bilingual financial modeling. It provides unrestricted access to diverse model variants, a substantial compilation of diverse cross-lingual and multi-modal instruction data, and an evaluation benchmark with expert annotations, comprising 10 NLP tasks, 20 bilingual specific tasks, totaling 95k datasets. Our thorough evaluation emphasizes the advantages of incorporating these bilingual datasets, especially in translation tasks and utilizing original English data, enhancing both linguistic flexibility and analytical acuity in financial contexts. Notably, ICE-INTENT distinguishes itself by showcasing significant enhancements over conventional LLMs and existing financial LLMs in bilingual milieus, underscoring the profound impact of robust bilingual data on the accuracy and efficacy of financial NLP.
{CJK}
UTF8gkai
@@ 1 Introduction
Beyond the efforts in English-centric models, Chinese financial LLMs (FinLLMs) have also made notable advances. XuanYuan2.0 (Zhang et al., 2023d) exemplifies this trend, focusing on the intricacies of Chinese financial language. This has been followed by models such as DISC-FinLLM (Chen et al., 2023) and CFGPT (Li et al., 2023a), which further the capabilities in financial analysis and reasoning. Recently, PanGu-π (Wang et al., 2023) enhances a pretrained model for better performance in financial examinations. Moreover, certain business-oriented LLMs (Wang, 2023) like TigerBot and LightGPT remain closed-source.
@@
BloombegGPT (Wu et al., 2023) 50B en 0 5 0 5 0 $\times$× $\times$× $\checkmark$ 03/30/23
InvestLM (Yang et al., 2023c) 65B en 0 9 0 9 0 $\times$× $\times$× $\checkmark$ 09/14/23
FinGPT (Wu et al., 2023) 7/13B en 0 6 0 6 0 $\times$× $\times$× $\checkmark$ 11/10/23
XuanYuan2.0 (Zhang et al., 2023d) 176B zh - 0 0 1 0 $\times$× $\times$× $\checkmark$ 05/19/23
FinMA (Xie et al., 2023b) 7/13B en 0 8 0 8 0 $\checkmark$ $\checkmark$ $\times$× 06/01/23
CFGPT (Li et al., 2023a) 7B zh 6 0 6 0 0 $\checkmark$ $\times$× $\checkmark$ 09/01/23
DISC-FinLLM (Chen et al., 2023) 13B zh 8 0 8 0 0 $\checkmark$ $\times$× $\checkmark$ 10/24/23
PanGu-π  (Wang et al., 2023) 1/7B zh - 0 4 0 0 $\times$× $\checkmark$ $\checkmark$ 12/27/23
ICE-INTENT 7B zh, en 12 6 12 6 3 $\checkmark$ $\checkmark$ $\times$× 12/10/23
Table 1: Overview of various FinLLMs. We outline their model size, language proficiency, instruction and evaluation counts in Chinese [zh], English [en] and data-out-of-field [dof] (not for training) financial specific tasks, open source (Yes $\checkmark$ and No $\times$×), data modality (text-only $\checkmark$ or not $\times$×) and release date. The task counts for all models are based on specific-task types detailed in Appendix Table 7, with "-" indicating unspecified information in their sources.
@@
In the creation of ICE-INTENT, we meticulously gathered 40 datasets, which consisted of 1,185,076 raw data, 603,940 instruction fine-tuning data, and 95,091 evaluation data, strategically covering a wide range of financial tasks. This ensemble, as detailed in Table 1, comprises 18 specific task on Chinese-English bilingual instruction datasets. This diverse array not only underscores our model’s commitment to linguistic breadth but also to task-specific depth, ensuring ICE-INTENT’s proficiency across various financial scenarios  [^2] Detailed examples are shown in Appendix Figure 8-9. Complementing this, ICE-FLARE, our cross-lingual evaluation benchmark, incorporates 21 evaluation specific tasks—12 in Chinese and 9 in English. Particularly, we innovatively introduced data-out-of-field specific task into ICE-FLARE. This robust framework highlights the unique strength of our approach in providing comprehensive cross-lingual consistency and generalization assessments, crucial for bilingual FinLLMs.
@@ 2 Related Work
Financial Dataset Resource: Financial pre-trained language models (FinPLMs) previously focused on single-task data but now most FinLLMs prioritize specific-task data collection. For instance, BloombergGPT (Wu et al., 2023) integrates general and financial text for training large 500B LLMs. FinLLMs like FinGPT (Yang et al., 2023b; Zhang et al., 2023a) curate English instruction datasets for classification, extraction, and generation NLP tasks. PIXIU  (Xie et al., 2023b) provide open-source, multi-task and multi-modal data covering diverse financial scenarios. BBT-Fin (Lu et al., 2023) Corpus aggregates Chinese financial data including corporate announcements, research reports, social media, and more, while XuanYuan (轩辕) 2.0 (Zhang et al., 2023d) offers a hybrid-tuning dataset. Cornucopia’s (Yu, 2023) 26M instruction dataset covers application areas like insurance and stocks. Most fine-tuning datasets are tailored for monolingual tasks, limiting coverage and open sourcing.
@@ 3.2.1 Chinese Financial Datasets
For ZH-CLS, our dataset covers essential tasks such as sentiment analysis (FinSA), semantic matching (FinSM), news classification (FinNC), negative judgment (FinNJ), answer selection (FinAS) and relationship extraction (FinRE). 1) Sentiment Analysis. It is the process of analyzing and labeling the sentiment expressed in financial texts as "Negative(消极)","Neutral(中性)" and "Negative(消极)" (Sohangir et al., 2018). We used the FE (Lu et al., 2023) dataset sourced from Stockbar and Xueqiu websites. We also utilized the StockB   [^3] https://huggingface.co/datasets/kuroneko5943/stock11 dataset, which focuses on the finance in among data from ten domains. 2) Semantic Matching. It determines if financial texts or sentences are semantically similar ("是(YES)" or "否(NO)"), using two datasets: the Bank Question Corpus (BQC)  (Chen et al., 2018) from Chinese bank customer logs and the Ant Financial Question Matching Corpus (AFQMC)  (Xu et al., 2020) from the Alipay competition, both utilizing binary classification for match identification. 3) Negative Judgment. It focuses on identifying finance and economics-related negativity, distinct from general financial sentiment analysis. Leveraging the FinNSP dataset (Lu et al., 2023) sourced from social media, it uses "是(YES)" to label data containing negative information and "否(NO)" for those without. 4) News Classification. It is to categorize news into 2 or multiple-class label, such as "股票市场(stock market)", "外汇(foreign exchange)", "宏观经济(macroeconomics)", and more. We use the FinNL  (Lu et al., 2023) dataset, sourced from financial news websites like Sina Finance, Tencent Finance, Phoenix Finance, 36Kr, and Huxiu. It includes binary (NL) and multi-class (NL2) labeled data. 5) Answer Selection. FinEval  (Zhang et al., 2023b) and and AntGroup evaluates financial reasoning by selecting the correct option ("A", "B", "C", "D) in a financial context. The dataset   [^4] https://github.com/alipay/financial_evaluation_dataset includes 4,661 questions covering finance, economy, accounting, and certifications. We’ve compiled data for ten key finance subjects as FinEvalF. 6) Relationship Extraction. It aims to provide the types of connections, dependencies, and associations between entities. We gathered the FinRE dataset (Lu et al., 2023) and named it RE, which includes financial news articles along with head-tail entity pairs, covering 44 relationship categories (e.g., "合作(cooperation)", "持股(shareholding)", etc.), including the "未知(unknown)" relation.
@@ 3.2.2 English Financial Datasets
For EN-EXT, we include the NER task. We use the CoNLL-2003 dataset (Alvarado et al., 2015) named as NER including sentences from public financial agreements via U.S. SEC filings. It includes four annotated entity types: Location ("LOC"), Organisation ("ORG"), Person ("PER"), and Miscellaneous ("MISC").
@@
Additionally, we introduce EN-OFT including NER, text summarization (FinTS) and hawkish-dovish classification (FinDC) task. 1) Named Entity Recognition. In NER task, we utilize the FINER-ORD  (Shah et al., 2023b) dataset, which provides labels for entities such as "PER", "LOC", and "ORG". 2) Text Summarization. For this task, we use two datasets: ECTSUM (Mukherjee et al., 2022) and EDTSUM (Zhou et al., 2021). ECTSUM involves extracting key sentences from texts, while EDTSUM focuses on generating fitting news headlines through abstractive summarization. 3) Hawkish-dovish Classification. FinDC aims to classify sentences from monetary policy texts into a "hawkish" or "dovish" stance, unlike FinSA, using the FOMC  (Shah et al., 2023a) dataset.
@@ 4 ICE-FIND: Bilingual Financial Instruction Data
Base on the raw datasets, we further construct our bIlingual Chinese-English Financial InstructioN Dataset (ICE-FIND), covering 13 tasks and 36 datasets with 604k data samples in total, as detailed in Appendix Table 7. This dataset compilation combined raw data with expertly devised prompts, generating 30-40 unique prompts for each dataset. However, it is important to note that expert-authored prompts may not always accurately reflect the quality of the generated results. We assessed these prompts through human evaluations  [^6] Examples and thier translations are shown in Appendix Figure 8 focusing on accuracy, naturalness, and informativeness using a 1-3 scale, to identify the most effective ones. These prompts were tested on various LLMs such as Bing Chat, ChatGLM using the ChatALL platform Sun (2024), showing satisfactory performance in most cases.
@@
To handle prompt variations across LLMs, we adopte two strategies. For English tasks, we match each data with all prompts except for FinQA and ConvFinQA. For Chinese tasks, we randomly select a prompt for each data due to the larger scale of the raw data. Finally, we convert the raw datasets into instruction-tuning samples by by restructuring expert prompts  [^7] Examples are presented in Appendix Table 10, input texts, and responses. The instruction data was structured in JSON format (Figure 3). {data_id} represents the ID of each data, {prompt} is the task-specific prompt, {conext} refers to the input data, {Your answer} denotes the expected answer, and {Your label} represents the truly answer.
@@ 5 ICE-INTERN: Bilingual Financial Large Language Model
Table 2: The details of specific task, evaluation metric, language type, dataset name and test size in ICE-FLARE. "Cover" indicates whether (Yes $\checkmark$ and No $\times$×) the public evaluation benchmark ([1]FLARE, [2]CFBenchmark, [3]FinCUGE, [4]FLUE, [5]Fineval, [6]CGCE) cover the specific task of ICE-FLARE.
@@
Sentiment Analysis FinSA F1 Accuracy zh FE 2,020 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$× $\times$×
@@
F1 Accuracy zh Corpus AFQMC 10,000 4,316 $\times$× $\times$× $\times$× $\times$× $\times$× $\times$×
@@
F1 Accuracy zh NL NL2 884 884 $\times$× $\checkmark$ $\checkmark$ $\times$× $\times$× $\times$×
@@
F1 Accuracy zh NSP 500 $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\times$×
@@
F1 Accuracy zh FinevalF 222 $\times$× $\times$× $\checkmark$ $\times$× $\checkmark$ $\times$×
Stock Prediction FinSP F1 Accuracy MCC zh StcokA 1,477 $\checkmark$ $\times$× $\times$× $\times$× $\times$× $\times$×
@@
F1 Accuracy zh RE 1,489 $\times$× $\times$× $\checkmark$ $\times$× $\times$× $\times$×
@@
Avg F1 zh en CHeadlines Headlines 2,051 20,547 $\checkmark$ $\times$× $\times$× $\checkmark$ $\times$× $\times$×
@@
F1 Accuracy MCC en German Australian 200 139 $\times$× $\times$× $\times$× $\times$× $\times$× $\times$×
@@
F1 Accuracy en FOMC 496 $\times$× $\times$× $\times$× $\times$× $\times$× $\times$×
Question Answering FinQA EM Accuracy zh QA 2,469 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$× $\checkmark$
@@
F1 Entity F1 zh en en CNER NER FINER-ORD 337 98 1,075 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$× $\times$×
Event Detection FinED F1 Recall Precision zh 19CCKS 2,936 $\times$× $\checkmark$ $\times$× $\times$× $\times$× $\times$×
@@
Rouge BERTScore BARTScore zh en en NA ECTSUM EDTSUM 3,600 495 2000 $\times$× $\checkmark$ $\checkmark$ $\times$× $\times$× $\times$×
Open Source $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$×
@@
Upon evaluating performance of various backbone LLMs on ICE-FLARE, we explore InternLM-7B to bulid ICE-INTERN-7B with the ICE-FIND. To assess the impact of different data types on fine-tuning ICE-INTERN, we gradually incorporate the DLC, DLE, DTT and DTE datasets during the fine-tuning process to create variants of ICE-INTERN named ICE-INTERN-dlc-7B, ICE-INTERN-dle-7B, ICE-INTERN-dtt-7B, and ICE-INTERN-full-7B, respectively. We use the same set of fine-tuning hyper-parameters for all models, which are described below. To improve the model’s capability in handling long text, we employ an advanced fine-tuning technique called QLoRA Hu et al. (2021) to enhance the training process, leading to optimized memory usage and faster training times. Specifically, the training sequences are uniformly cut to a length of 2048 tokens. Then, we use AdamW optimizer with an initial learning rate 5e-5 on weight decay 1e-5 for all parameters. Additionally, the warmup steps are set to 1% of the total training steps. Finally, we fine-tune model total 1 epoch in batch size of 24 on 8 A100 40GB GPUs. Due to the growing volume of fine-tuning data and constraints on our computational resources, we have opted not to proceed with additional fine-tuning of ICE-INTERN using larger parameters.
6 ICE-FLARE: Bilingual Financial Evaluation Benchmark
Table 3: The performance of different 6B-7B general LLMs, baseline FinLLMs and variants of ICE-INTERN on the ICE-FLARE benchmark. All performance results are based on the average results of the NLP tasks including classification (CLS), extraction (EXT), reasoning (REA), prediction (PRE), generation (GEN), translation (TRA), and data-out-of-field (DOF) serving as the final evaluation criteria of ICE-FLARE benchmark. Results with "__" indicate the best results among models excluding ChatGPT.
Model ZH-CLS ZH-PRE ZH-EXT ZH-GEN ZH-TRA ZH.Avg EN-CLS EN-PRE EN-EXT EN-REA EN-OFT EN.Avg Bi.Avg CLS.Avg PRE.Avg EXT.Avg
Bloomz-7B1 0.170 0.131 0.038 -1.650 0.240 -0.214 0.406 0.238 0.000 0.005 -1.238 -0.118 -0.166 0.288 0.185 0.019
Baichuan-7B 0.202 0.000 0.005 -2.097 0.241 -0.330 0.264 0.248 0.000 0.000 -1.246 -0.147 -0.238 0.233 0.124 0.003
Baichuan2-7B 0.005 0.000 0.001 -2.262 0.179 -0.415 0.173 0.226 0.000 0.000 -1.307 -0.182 -0.299 0.089 0.113 0.001
LLaMa2-7B 0.205 0.000 0.001 -1.262 0.294 -0.152 0.271 0.243 0.012 0.000 -1.214 -0.138 -0.145 0.238 0.122 0.007
ChatGLM2-6B 0.036 0.000 0.005 -2.265 0.179 -0.409 0.205 0.225 0.000 0.000 -1.298 -0.174 -0.291 0.121 0.113 0.003
ChatGLM3-6B 0.391 0.220 0.013 -1.031 0.311 -0.019 0.434 0.262 0.248 0.000 -0.833 0.022 0.002 0.413 0.241 0.131
Qwen-7B 0.327 0.154 0.001 -1.185 0.295 -0.082 0.454 0.226 0.012 0.000 -1.173 -0.096 -0.089 0.391 0.190 0.007
ChatGPT 0.432 0.422 0.039 -1.693 0.311 -0.098 0.430 0.262 0.770 0.590 -0.719 0.267 0.084 0.431 0.342 0.405
GPT-4 0.549 0.383 0.098 -1.577 0.266 -0.056 0.618 0.283 0.830 0.695 -0.640 0.357 0.151 0.584 0.333 0.464
InternLM-7B 0.277 0.003 0.001 -1.119 0.272 -0.113 0.301 0.236 0.000 0.000 -1.251 -0.143 -0.128 0.289 0.120 0.001
CFGPT-sft-7B 0.211 0.050 0.011 -2.017 0.272 -0.295 0.242 0.224 0.000 0.000 -1.265 -0.160 -0.227 0.227 0.137 0.006
DISC-FinLLM-13B 0.290 0.315 0.022 -1.858 0.290 -0.188 0.446 0.236 0.039 0.002 -1.158 -0.087 -0.138 0.368 0.276 0.031
FinMA 0.229 0.000 0.016 -2.009 0.250 -0.303 0.371 0.233 0.392 0.000 -1.187 -0.038 -0.171 0.300 0.117 0.204
ICE-INTERN-dlc-7B 0.598 0.340 0.020 -1.762 0.337 -0.093 0.331 0.282 0.000 0.000 -1.227 -0.123 -0.108 0.465 0.311 0.010
ICE-INTERN-dle-7B 0.607 0.637 0.569 -1.024 0.327 0.223 0.334 0.275 0.006 0.000 -1.203 -0.118 0.053 0.471 0.456 0.288
ICE-INTERN-dtt-7B 0.607 0.630 0.566 -1.024 0.448 0.245 0.357 0.268 0.000 0.000 -1.182 -0.111 0.067 0.482 0.449 0.283
ICE-INTERN-full-7B 0.631 0.617 0.567 -0.972 0.409 0.250 0.494 0.250 0.362 0.000 -1.193 -0.017 0.117 0.563 0.434 0.465
Based on ICE-FIND, we design our bilingual multi-task evaluation benchmark, ICE-FLARE. The existing benchmarks ([1] FLARE Xie et al. (2023b), [2] CFBenchmark Lei et al. (2023), [3] FinCUGE Chen et al. (2023), [4] FLUE Shah et al. (2022), [5] Fineval Zhang et al. (2023b), [6] CGCE Zhang et al. (2023c)) compared for ICE-FLARE are listed in Table 2. Unfortunately, they are all designed exclusively for either Chinese or English. ICE-FLARE has highlighted its superiority over FLUE in English through FLARE benchmark Xie et al. (2023b). From Table 2, it is evident that ICE-FLARE introduce bilingual capabilities, significantly surpasses these benchmarks on task diversity and data scalability. Despite FinCUGE having a higher number of financial evaluation tasks, the difference in dataset size compared to ICE-FLARE is nearly 8 times. These public benchmarks have provided valuable references and inspiration for constructing the test set, making ICE-FLARE more comprehensive in terms of task coverage, metric diversity, and sample quantity. However, they still face challenges in evaluating the bilingual capabilities in Chinese-English and severely lack DOF tasks (e.g., FinDC) to test the robustness of the models. In conclusion, ICE-FLARE, which offers a bilingual evaluation benchmark and a more comprehensive assessment of diversity tasks, is necessary for improving performance and linguistic capabilities of FinLLMs.